VIBRONIC STRUCTURS OF THE MgNC Ã ^{2}Π AND X̃ ^{2}Σ^{+} STATES

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated MgNC in supersonic free jet expansions, and measured the laser induced fluorescence excitation spectra of the vibronic bands of the Ã ^{2}Π - X̃ ^{2}Σ^{+} transition. We have found some new features in the spectra, though part of the spectra were already reported by Wright and Miller^{a}. The new features are consisting of the ν_{2} bending vibronic bands. From the rotational analyses of some of the vibronic bands, we have obtained information on the vibronic structure of the Ã ^{2}Π state of MgNC. The assignments of the ν_{2} vibronic bands were carried out utilizing the vibrational analyses of the dispersed fluorescence spectra obtained by the excitation of the ν_{2} vibronic bands. On the basis of the vibrational structures of the dispersed fluorescence spectra which are not only those of the ν_{2} vibronic bands but also the spectra obtained by the excitation of the other vibronic bands reported by us^{b}, the vibronic structure of the X̃ ^{2}Σ^{+} state of MgNC has been analysed. The vibronic structures of the Ã ^{2}Π and X̃ ^{2}Σ^{+} states of MgNC will be discussed comparing with the computational results by Hirano group^{c}.

aR. R. Wright and T. A. Miller, J. Mol. Spectrosc. \underline{194}, 219 (1999).

bM. Fukushima and T. Ishiwata, J. Mol. Spectrosc. \underline{216}, 159 (2002).

cT. E. Odaka, T. Taketsugu, T. Hirano and U. Nagashima, J. Chem. Phys. \underline{115}, 1349 (2001); T. E. Odaka, T. Hirano and P. Jansen, J. Mol. Spectrosc. \underline{211}, 147 (2002); T. Hirano, K. Ishii T. E. Odaka and P. Jansen, J. Mol. Spectrosc. \underline{215}, 42 (2002); T. E. Odaka, T. Hirano and P. Jansen, J. Mol. Spectrosc. \underline{216}, 379 (2002).